OpenStudy (anonymous):

How to find the inverse of y=1/x+1

OpenStudy (anonymous):

is this \[y = \frac{ 1 }{ x + 1 }\] or \[y = \frac{ 1 }{ x } + 1\] ?

OpenStudy (anonymous):

the first one

OpenStudy (anonymous):

to find the inverse, swap the variables: \[x = \frac{ 1 }{ y + 1 }\] then solve for y

OpenStudy (anonymous):

i know, but would I divide by 1 then take away the positive one from each side?

OpenStudy (anonymous):

well first you would interchange (y + 1) with x to get \[y + 1 = \frac{ 1 }{ x }\]

OpenStudy (anonymous):

oh ok. thanks